OpenStudy (anonymous):

A box has 8 beads of the same size, but all are different colors. Tamy draws a bead randomly from the box, notes its color, and then puts the bead back in the box. She repeats this 3 times. What is the probability that Tamy would pick a red bead on the first draw, then a white bead, and finally a red bead again? Answers are 1/4096 1/512 1/64 and 1/8 its not i1/8 cause I tried it and got it wrong so I though it would be 1/64? :)

OpenStudy (anonymous):

Hell o again

OpenStudy (kropot72):

The probability of drawing any one of the colors is 1/8. The results of each draw are independent. Therefore the required probability is \[\large \frac{1}{8}\times\frac{1}{8}\times\frac{1}{8}=you\ can\ calculate\]

OpenStudy (anonymous):

hi. so it's 1/8 for each draw since the beads are being replaced. Then you need to multiply them to get the probability \[\frac{ 1 }{ 8 }\times \frac{ 1 }{ 8 }\times \frac{ 1 }{ 8 }\]

OpenStudy (anonymous):

oh I got it thank you guys so much